the graph that represents a 1-1 function?
Answer by rapaljer(4671)

Any linear graph except for vertical and horizontal lines would do.
y=mx+b

Also, a cubic equation like y=x^3, y =x^3 + 2, etc.

Dr. Robert J. Rapalje,